OpenStudy (anonymous):

On a vacation, Bridget and her family traveled 225 miles at 67 mph. How long did the trip take? (Round to the nearest hundredth) 3.5 hours 3.25 hours 3.36 hours 3.30 hours

OpenStudy (texaschic101):

s = d/t but since you are looking for time, let me change the equation... t = d/s (time = distance/speed t = 225/67 = 3.358 rounds to 3.36

OpenStudy (anonymous):

so 3.36 is the answer

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

thanks